consider the curve described by the equation:

4x2 - 3xy + y2 = 14

at any given point on this curve, we have

dy/dx = -8x + 3y / -3x + 2y

your task is to find the points on the curve where the tangent line is parallel to the line y = x

What is the y-coordinate of the leftmost point on the curve where the tangent line is parallel to the lone y=x
Answer #1

This is the proper solution of your question.

Thank you.

4x² - 3 xy + y2 = 14 dy - - 8x+34 -3x+2y L 34 - 8 = 24 - 3x y = -3x+8x= Sx Jy sa 4x²-3x (Sx) + (5x)² = 14 4x²-15 x² + 25 x² =